# Bridging Earth and Space: A Survey on HAPS for Non-Terrestrial Networks

### v1

@@ I Introduction
The main objective of this survey is to present a comprehensive overview of the state-of-the-art in HAPS networks, examining key applications, enabling technologies, and future research directions. We first outline the diverse roles of HAPS across several domains, including connectivity enhancement, network infrastructure and management, IoT and emerging services, advanced control and special applications, as well as energy and sustainability. The survey further examines TN-NTN integration, emphasizing the role of HAPS as a bridge between terrestrial and non-terrestrial systems. It outlines platform and terminal types within this framework, reviews TN-NTN architectures, and explores emerging frequency bands such as millimetre wave (mmWave) and centimeter wave (cmWave), underscoring their importance for enhanced connectivity. The survey then reviews key HAPS technical enablers, including channel modeling, user equipment (UE) association and resource management, mobility and multi-link connectivity, massive multiple-input multiple-output (MIMO) and machine learning (ML) applications, interference control, and energy efficiency. It also addresses deployment and operational challenges, highlighting areas needing further research to fully realize HAPS potential. By covering these topics, the survey offers an in-depth understanding of technological advancements, design considerations, and open challenges in HAPS-based fifth-generation (5G) and 6G solutions, serving as a valuable resource for researchers and engineers in TN-NTN integration.
I-A Review of Previous NTN-Based Surveys
@@
I-A1 General NTN Surveys
Several surveys analyze NTNs, but few provide an in-depth examination of HAPS. The study in [4] explores UAVs in future communication systems, covering use cases, architectures, and challenges, including sub-6 GHz and mmWave TN optimization and UAV-to-UAV performance. The work in [5] reviews aerial connectivity for airplanes, vertical take-off and landing (VTOL) vehicles, and UAVs, analyzing use cases, requirements, and providers such as satellites and HAPS, though its coverage of HAPS is limited and lacks TN-NTN integration. Study [1] examines TN-NTN integration in 5G and 6G, focusing on mmWave, IoT/ Mobile Edge Computing (MEC) integration, and ML, but emphasizes LEO satellites and UAVs over HAPS. Similarly, [6] and [7] discuss LEO-focused TN-NTN integration, spectrum, MIMO modeling, mobility management, third generation partnership project (3GPP) standardization, and NTN evolution toward 6G, but only briefly mention HAPS. Overall, these surveys provide limited insight into HAPS as a central component of NTN networks.
I-A2 AI-Focused Surveys
[…]
AI is essential for NTNs and TN-NTN integration, enabling intelligent resource management, optimization, and automation across network layers. Several surveys examine AI in NTNs, focusing on key technologies, ML-based algorithms, and performance improvements. Study [8] investigates ML for UAV applications, including image processing, ground sensor localization, real-time control, scheduling, and trajectory planning. Reference [9] reviews 6G aerospace networks, analyzing satellite configurations, architectures, protocols, and AI-driven resource management, while highlighting research gaps. Work [10] explores ML techniques for NTN challenges in 6G, addressing privacy risks and optimization via both, deep learning and deep reinforcement learning (DRL). Study [11] analyzes AI-based strategies for TN-NTN integration, convergence drivers, key applications, and architectures. Despite this coverage, these surveys primarily focus on satellites and UAVs, with limited discussion of HAPS.
I-A3 HAPS-Related Surveys
The survey in [3] provides a comprehensive tutorial on HAPS and low-altitude platform stations, highlighting their roles as aerial base stations and cellular-connected UE. It examines challenges such as three-dimensional (3D) deployment, channel modeling, and energy efficiency, and reviews analytical frameworks including optimization theory and ML for UAV-related issues, but does not address TN-NTN integration. Study [12] reviews LEO satellite and HAPS communications, covering system design, networking, medium access control, air interface, testbeds, and NTN integration, yet focuses primarily on satellites. Work [13] considers NTNs for 6G, simulating TN-LEO and TN-LEO-HAPS integration, but treats HAPS only as relays. The survey in [2] focuses on HAPS in mega-constellation NTNs, exploring super macro base stations, regulatory aspects, channel modeling, resource management, handoff, swarm control, AI applications, and emerging technologies such as reconfigurable intelligent surfaces and faster-than-Nyquist signaling, yet lacks insights on TN-NTN interoperability. Finally, [14] examines HAPS for coverage extension and capacity enhancement in TN-HAPS coexistence, discussing intelligent resource management, interference mitigation, spatial optimization, and past projects, but does not address advanced applications like IoT support, AI-driven optimization, or fully integrated TN-NTN architectures.
Among the most relevant surveys [3, 12, 13, 2, 14], none offer a comprehensive, up-to-date analysis of HAPS covering both technological advancements and TN-NTN integration. Studies [3, 12] provide general NTN overviews with limited HAPS discussion, focusing on use cases, architectures and capabilities. While [13] addresses HAPS in NTN-enabled 6G, it emphasizes relay-based integration rather the more powerful regenerative-based one based on independent network nodes in the sky. Survey [2] reviews HAPS mega-constellations, regulations, and enabling technologies, yet offers limited insight on NTN deployment and TN-NTN interoperability. Work [14] considers coverage extension and terrestrial coexistence, but does not cover broader TN-NTN integration, advanced use cases, or emerging AI-related technologies. To address these gaps, this survey provides a holistic, forward-looking review of HAPS, including TN-NTN integration, enabling technologies, real-world deployment challenges, and future research directions.
I-B Survey Structure
The rest of this survey is organized as follows Section II compares HAPS with other NTN nodes. Section III highlights the most promising use cases for HAPS. Section IV explores scenarios and architectures for HAPS deployment. Section V examines the key technical enablers for HAPS. Section VI discusses the current challenges related to HAPS communications. Section VII concludes the survey with findings and insights for future research directions.
List of Abbreviations
@@ I-B Survey Structure
HARQ
hybrid automatic repeat request
IoRT
Internet of remote things
@@
RIS
reconfigurable intelligent surface
@@
II HAPS Comparison with Terrestrial Networks and Other NTN Nodes
Global connectivity, traditionally supported by TNs and satellites, still faces significant gaps. Despite mobile broadband traffic doubling from 419 exabytes (EB) in 2019 to 913 EB in 2022, about 33% of the world’s population remains offline [15]. This digital divide highlights the need for innovative solutions. HAPS, operating in the stratosphere at 20–50 km, offer a promising approach. Acting independently or as intermediaries between TNs and other NTN systems such as LEO satellites and UAVs, HAPS provide wide-area coverage, low latency, and operational flexibility. These features enable them to bridge connectivity gaps while complementing existing networks. This section examines HAPS capabilities compared with TNs (Section II-A), LEO satellites (Section II-B), and UAVs (Section II-C), highlighting their unique advantages. Table I summarizes the distinct characteristics of HAPS relative to TNs and other NTN nodes in terms of altitude, performance, operational flexibility, and cost.
Table I: Comparison of HAPS with TNs, LEO satellites, and UAVs.
@@ II HAPS Comparison with Terrestrial Networks and Other NTN Nodes
II-A Comparison with Terrestrial Networks
@@
TNs serve as the backbone of modern communication systems, providing high-capacity, low-latency connectivity in densely populated areas. However, they are constrained by geographical and infrastructural limitations. Rugged terrains, remote regions, and disaster-stricken areas often lack adequate terrestrial infrastructure, leaving significant portions of the population unconnected. Furthermore, the deployment and maintenance of TNs in such regions involve high costs, making them impractical for universal connectivity. HAPS overcome these limitations by offering wide-area coverage without reliance on extensive ground infrastructure. They can provide reliable LoS connectivity across urban, rural, and remote areas, bridging gaps where terrestrial solutions are infeasible. Their rapid deployment capability also makes them ideal for emergency scenarios, while operational costs in remote areas are often lower than constructing terrestrial networks.
II-B Comparison with LEO Satellites
LEO satellites provide essential global coverage, particularly in remote and oceanic regions. However, their high operational altitudes (500 to 2,000 kilometers) result in increased signal attenuation and latency. Moreover, their large coverage areas imply that the bandwidth is shared among numerous UEs, leading to lower per-UE data rates. These constraints make LEO satellites less suited for high-bandwidth or latency-sensitive applications. In contrast, HAPS are positioned closer to Earth, which significantly reduces latency and enables a better spatial reuse, resulting in an enhanced bandwidth availability per UE and higher data rates. Additionally, HAPS are more operationally flexible than LEO satellites. While satellites require complex and costly launches, HAPS can be repositioned and upgraded with minimal disruption, making them adaptable to evolving connectivity demands.
II-C Comparison with UAVs
UAVs are highly versatile, providing localized and flexible connectivity for hot-spot tactical operations and disaster scenarios. However, their limited flight duration due to battery power and energy constraints, payload capacity, and frequent maintenance requirements hinder scalability in large or long-term deployments. In contrast, HAPS offer extended operational periods, lasting weeks or even months, powered by solar energy and autonomous systems. Their quasi-stationary positions enable them to serve as stable, high-capacity communication hubs, bridging the gap between the localized capabilities of UAVs and the global reach of satellites.
III HAPS Use Cases

### v2

@@ I Introduction
The main objective of this survey is to present a comprehensive overview of the state-of-the-art in HAPS networks, examining key applications, enabling technologies, and future research directions. We first outline the diverse roles of HAPS across several domains, including connectivity enhancement, network infrastructure and management, IoT and emerging services, advanced control and special applications, as well as energy and sustainability. The survey further examines TN-NTN integration, emphasizing the role of HAPS as a bridge between terrestrial and non-terrestrial systems. It outlines platform and terminal types within this framework, reviews TN-NTN architectures, and explores emerging frequency bands such as millimetre wave (mmWave) and centimeter wave (cmWave), underscoring their importance for enhanced connectivity. The survey then reviews key HAPS technical enablers, including channel modeling, resource management, user equipment (UE) association, multi-link connectivity, mobility management and handover, antenna technologies, interference control, energy efficiency, machine learning (ML) applications, and cybersecurity. It also addresses deployment and operational challenges, highlighting areas needing further research to fully realize HAPS potential. By covering these topics, the survey offers an in-depth understanding of technological advancements, design considerations, and open challenges in HAPS-based fifth-generation (5G) and 6G solutions, serving as a valuable resource for researchers and engineers in TN-NTN integration.
The literature for this survey was sourced through a manual, systematic search across key academic databases, including IEEE Xplore, Scopus, and Google Scholar, resulting in the selection of over 200 peer-reviewed studies. The search prioritized publications from the last decade (2015–2025) to capture recent advancements in HAPS technologies, with keywords such as "HAPS," "stratospheric platforms," "6G networks," "non-terrestrial networks" etc., and related terms in telecommunications and radiophysics. An initial screening was conducted based on titles and abstracts to identify potentially relevant works, followed by a full-text review to apply inclusion criteria focused on topic relevance to HAPS/NTN-based system models, and TN-NTN integration; methodological rigor, including empirical validation or theoretical consistency with standards like International Telecommunication Union Radiocommunication Sector (ITU-R) recommendations; and comprehensiveness in addressing challenges. Exclusions were made for a few non-peer-reviewed materials (arXiv preprints, press-releases and reports), relevant studies predating 2015, or those lacking substantive contributions, but ensuring a comprehensive yet focused corpus that supports the survey’s objectives.
The rest of this survey is organized as follows (see Fig. 1): Section II reviews the most relevant prior surveys including general studies in the area of non-terrestrial networks, studies focused on AI and ML applications and those specifically considering HAPS. Section III compares HAPS with other NTN nodes. Section IV highlights the most promising use cases for HAPS. Section V explores scenarios and architectures for HAPS deployment. Section VI examines the key technical enablers for HAPS. Section VII discusses the current challenges related to HAPS communications. Section VIII concludes the survey with findings and insights for future research directions.
Figure 1: Survey structure
II Review of Previous NTN-Based Surveys
@@
II-A General NTN Surveys
Several surveys analyzed NTNs, but most provided no in-depth examination of HAPS. Baltaci et al. [36] reviewed aerial connectivity for airplanes, vertical take-off and landing vehicles, and UAVs, providing a detailed analysis of use cases, requirements, and network architectures. It also examined aerial connectivity providers, including satellites and HAPS, as well as communication challenges. However, its coverage of HAPS is limited, and lacks examination of TN-NTN integration. Geraci et al. [73] explored the role of UAVs in future communication systems, covering use cases, network architectures, and research challenges. It reviewed potential optimization opportunities in sub-6 GHz and mmWave TNs to provide service to UAVs. Additionally, it included a performance analysis of UAV-to-UAV communications.
Azari et al. [32] studied TN-NTN integration in 5G and 6G networks, focusing on enabling technologies such as mmWave, IoT/ Mobile Edge Computing (MEC) integration, and ML. It also highlighted outcomes from field trials. However, it mostly emphasized LEO satellites and UAVs, rather than HAPS. TN-NTN integration was also examined by Geraci et al. [74], mostly focusing on LEO networks, spectrum challenges, multiple-input multiple-output (MIMO) modeling, and simulations to evaluate the performance of UAVs when served from either a massive MIMO-based TN or a LEO-satellite-based operator. The study also highlighted relevant third generation partnership project (3GPP) standardization efforts and solutions, such as three-dimensional (3D) radio access, mobility management, and network orchestration. Similarly, Rinaldi et al. [180] explored the evolution of NTN from second-generation to fourth-generation (4G), and discussed TN-NTN integration, mobility management, and 3GPP support for 5G new radio (NR). The survey also highlighted the role of NTN toward 6G for achieving global coverage and increased capacity.
Jamshed et al. [107] presented a comprehensive overview of NTN as a pivotal enabler for 6G networks, emphasizing their role in achieving seamless, intelligent, and sustainable global connectivity by integrating satellites, aerial platforms (UAVs, HAPS), and TNs to address the digital divide and support applications like IoT, navigation, disaster recovery, and interplanetary communication. The paper delved into enabling technologies, real-world applications illustrated through UAV/HAPS-assisted fronthaul/backhaul, security considerations, and future research directions, positioning NTN as essential part of resilient 6G ecosystems. Despite their comprehensive coverage of NTNs, these surveys only briefly mentioned HAPS as part of a broader NTN framework, and did not explore them in depth.
II-B AI-Focused Surveys
AI played an essential role in NTNs and TN-NTN integration, enabling intelligent resource management, optimization, and automation across various network layers. Several surveys analyzed the application of AI in NTNs, focusing on key technologies, ML-based algorithms, and network performance improvements. AI-based optimization strategies for TN-NTN integration were analyzed by Wang et al. [221], who discussed the potential of TN and satellite network integration to enhance global connectivity. This study provided an in-depth analysis of convergence drivers, network architectures, and key application areas. Kurunathan et al. [127] explored the use of ML for UAV-related applications, including image processing, ground sensors localization, real-time UAV control, scheduling and trajectory planning. Zhou, Sheng, Li and Han [252] reviewed 6G aerospace-integrated networks, analyzing architecture, satellite configurations, protocols, and key components, while also identifying future research challenges. It examined advanced technologies, with a focus on AI-driven resource management, and NTN research gaps. Iqbal et al. [101] also explored ML and AI techniques for addressing NTN-related challenges in 6G, highlighting privacy risks and optimization techniques such as deep learning, and deep reinforcement learning (DRL). Fontanesi et al. [68] provided a comprehensive overview of AI and ML applications in satellite communication (SatCom) and NTN, focusing on geostationary equatorial orbit (GEO), medium Earth orbit (MEO), and LEO systems, alongside with an extensive classification of the most important use cases, its related challenges and the conventional tools used to address them. Onboard vs. on-ground AI architectures were compared, emphasizing training paradigms (offline/online). Hardware solutions for ML deployment were reviewed, including radiation-tolerant chipsets. Long-term challenges and insights for future research concluded the study. Mahboob and Liu [146] categorized AI applications across current research directions covering physical (PHY) layer, data link layer and upper layer aspects such as computational offloading, network routing and traffic prediction. Existing works, research efforts, including ML and software defined radio testbeds, and practical challenges were reviewed, concluding with insights for future AI-driven NTN deployments.
Despite such a broad coverage, these surveys primarily focused on satellites and UAVs, with limited discussion of HAPS.
II-C HAPS-Related Surveys
Mozaffari et al. [162] provided a comprehensive tutorial on both HAPS and low-altitude platform stations in wireless communications, highlighting their role as aerial base stations and cellular-connected UEs. It examined key challenges such as 3D deployment, channel modeling, and energy efficiency, while exploring potential research directions. Additionally, it reviewed analytical frameworks, including optimization theory and ML, for addressing UAV-related issues. However, it did not consider TN-NTN integration, comparing instead UAV- and ground-based networks.
Kodheli et al. [124] reviewed advancements in LEO satellite and HAPS communications, capturing the state of the art in SatCom, and highlighting key developments such as new constellation architectures, on-board processing, and NTN integration. It systematically analyzed five critical areas: system design, air interface, medium access control, networking, as well as testbeds and prototyping. While the study covered HAPS, its primary focus remained on satellites, offering only a high-level assessment of HAPS without detailing their unique technological challenges. Giordani and Zorzi in [77] considered NTNs as a crucial enabler for 6G, enhancing connectivity through technologies such as mmWave and flexible beam steering. It complemented previous works by providing simulations that compared different integration scenarios, including TN-LEO, as well as TN, LEO, and HAPS integration. However, HAPS was considered only as a relay, without evaluating their potential as independent network nodes or fully-integrated NTN components. Karabulut Kurt et al. [116] took a more focused approach on HAPS, exploring their role in NTN communications, particularly in the context of mega-constellation systems for next-generation wireless networks. It examined key use cases, and focused on the potential of HAPS as super macro base stations, while also addressing regulatory considerations, alongside critical technologies such as channel modeling, resource management, handoff mechanisms, and swarm control. The survey also discussed AI applications, particularly in network topology management, resource allocation, and mobility optimization. Additionally, it reviewed emerging technologies such as reconfigurable intelligent surfaces for cost-effective deployment and faster-than-Nyquist signaling for spectral efficiency. While this survey provided a strong technological perspective, it lacked insights into TN-NTN interoperability and real-world deployment challenges. The role of HAPS in NTN was further explored by Arum, Grace, and Mitchell [30], who examined HAPS potential for coverage extension and capacity enhancement in TN-HAPS coexistence scenarios. The study focused on leveraging HAPS as an alternative to terrestrial systems, particularly for wireless communication in rural and remote areas where broadband access remained limited. It discussed techniques such as intelligent radio resource management, interference mitigation, and spatial optimization using array antennas, radio environment map, and device-to-device communication to maximize coverage while ensuring coexistence with TN. Additionally, the study highlighted past HAPS projects but did not explore new applications beyond coverage expansion, such as AI-driven optimization, network infrastructure and management, support for IoT, or TN-NTN integrated architectures. Abbasi et al. [11] outlined potential HAPS use cases and identified particular challenges which may arise due to integration of HAPS with conventional TNs and because of platform’s physical characteristics. The authors also investigated the usefulness of HAPS-enabled vertical heterogeneous networks through the estimating of cell-edge users’ capacity enhancement and the benefits of the cell-free communication system of aerial BSs with HAPS backhauling. Among the most relevant surveys to this work, i.e., [162, 124, 77, 116, 30, 11], none provided a comprehensive, up-to-date analysis of HAPS, covering both technological advancements and TN-NTN integration. The surveys in [162, 124] were still generalist NTN studies, offering only limited exploration of HAPS, focusing on use cases, technological capabilities, and network architectures, without detailing dedicated HAPS advancements. Although [77] considered HAPS in the NTN-enabled 6G ecosystem, it primarily focuses on relay-based NTN integration, without assessing HAPS as independent network nodes or their role in multi-layered communication systems. The survey [116] provided an extensive review of HAPS mega-constellations, regulatory aspects, and key enabling technologies, but offered little insight into TN-NTN interoperability, the opportunities and challenges it presents, and real-world deployment considerations. The work in [30] examined HAPS for coverage extension and terrestrial coexistence, but did not explore their broader integration into NTN architectures, advanced use cases beyond connectivity, or emerging technologies, especially those related to AI. While the paper [11] covered particular use cases and challenges it is more concise and less exhaustive in referencing. To fill these gaps, this survey provided a holistic and forward-looking review of HAPS, including TN-NTN integration, enabling technologies, real-world deployment challenges, and future research directions. Table I provided a comparative summary of this survey against the previous relevant literature.
Table I: Comparison of our survey against other HAPS-related surveys.
Survey Scope HAPS
focus
HAPS
use cases
HAPS
technologies
TN-NTN
integration
Challenges
[162] Examines HAPS and LAPS in wireless communications, identifying challenges and research directions but focusing on UAV-ground comparisons rather than TN-NTN integration. Secondary
[124] Reviews advancements in satellites and HAPS communication, highlighting key developments such as new constellations, on-board processing, and NTN integration, but focusing on satellite communications. Secondary
[77] Considers NTN for 6G, with simulations of TN-LEO-HAPS integration, but focusing on HAPS as a relay rather than independent nodes. Secondary
[116] Focuses on HAPS mega-constellations, regulatory aspects, and key enabling technologies, including AI and swarm control, but lacks TN-NTN integration insights. Primary
[30] Examines HAPS for coverage extension and coexistence with terrestrial networks, but does not explore broader use cases, advanced technologies, or TN-NTN integration. Primary
[11] Outline particular HAPS use cases and challenges in 6G vHetNets with some simulation validations. Primary
Ours Comprehensive review of HAPS, covering use cases, enabling technologies, TN-NTN integration, current HAPS-related projects, and future research directions. Primary
List of Frequently Used Abbreviations
@@ II-C HAPS-Related Surveys
GNSS
global navigation satellite system
GPS
global positioning system
@@
HO
handover
@@
ISTN
integrated space–terrestrial network
ITU
International Telecommunication Union
@@
RTT
round-trip time
@@
III HAPS Comparison with Terrestrial Networks and Other NTN Nodes
Global connectivity, traditionally supported by TNs and satellites, still faces significant gaps. Despite mobile broadband traffic doubling from 419 exabytes (EB) in 2019 to 913 EB in 2022, about 33% of the world’s population remains offline [105]. This digital divide highlights the need for innovative solutions. HAPS, operating in the stratosphere at 20–50 km, offer a promising approach. Acting independently or as intermediaries between TNs and other NTN systems such as LEO satellites and UAVs, HAPS provide wide-area coverage, low latency, and operational flexibility. These features enable them to bridge connectivity gaps while complementing existing networks. This section examines HAPS capabilities compared with TNs (Section III-A), LEO satellites (Section III-B), and UAVs (Section III-C), highlighting their unique advantages. Table II summarizes the distinct characteristics of HAPS relative to TNs and other NTN nodes in terms of altitude, performance, operational flexibility, and cost.
Table II: Comparison of HAPS with TNs, LEO satellites, and UAVs.
@@ III HAPS Comparison with Terrestrial Networks and Other NTN Nodes
III-A Comparison with Terrestrial Networks
@@
TNs serve as the backbone of modern communication systems, providing high-capacity, low-latency connectivity in densely populated areas. However, they are constrained by geographical and infrastructural limitations. Rugged terrains, remote regions, and disaster-stricken areas often lack adequate terrestrial infrastructure, leaving significant portions of the population unconnected. Furthermore, the deployment and maintenance of TNs in such regions involve high costs, making them impractical for universal connectivity.
Recent studies on networking have also emphasized these limits. For instance, previous research [59] illustrates the challenge of extending traditional terrestrial infrastructure into rural areas, while efforts on community cellular networks [88] and virtual coverage [92] reveal that operational complexity, cost, and power availability are the main obstacles to large-scale terrestrial deployment in underserved areas.
HAPS overcome these limitations by offering wide-area coverage without reliance on extensive ground infrastructure. They can provide reliable LoS connectivity across urban, rural, and remote areas, bridging gaps where terrestrial solutions are infeasible. Their rapid deployment capability also makes them ideal for emergency scenarios, while operational costs in remote areas are often lower than constructing terrestrial networks.
III-B Comparison with LEO Satellites
LEO satellites provide essential global coverage, particularly in remote and oceanic regions. However, their high operational altitudes (500 to 2,000 kilometers) result in increased signal attenuation and latency. Moreover, their large coverage areas imply that the available bandwidth is shared among a large number of UEs, leading to reduced per-UE data rates.
These system-level limitations are further supported by recent research. In LEO systems, downlink opportunities, contact windows, and ground-station coordination dominate end-to-end latency, according to work on distributed downlink architectures and constellation-scale traffic scheduling [217]. Direct-to-cell designs [137] highlight fundamental per-user capacity and resource-sharing limitations when large footprints serve massive user populations.These constraints make LEO satellites less suited for high-bandwidth or latency-sensitive applications. In contrast, HAPS are positioned significantly closer to Earth, which reduces latency and enables better spatial reuse, resulting in higher bandwidth availability per UE. Additionally, HAPS offer greater operational flexibility, as they can be repositioned and upgraded without the complexity and cost of satellite launches.
III-C Comparison with UAVs
UAVs are highly versatile, providing localized and flexible connectivity for hot-spot operations and disaster scenarios. However, their limited flight duration due to battery constraints, restricted payload capacity, and frequent maintenance requirements hinder scalability for large-area or long-term deployments. Recent system-level investigations also show these limitations. The efficiency of UAVs for short-term, localized coverage is demonstrated by research on drone-based communication relays [144] and integrated rural connectivity platforms [216], while also highlighting endurance and sustainability constraints that preclude persistent wide-area service.
In contrast, HAPS offer extended operational durations, lasting weeks or months, enabled by solar power and autonomous control. Their quasi-stationary positioning allows them to function as stable, high-capacity communication hubs, effectively bridging the gap between the localized flexibility of UAVs and the global reach of satellite systems.
III-D Quantitative and Structured Comparison Across NTN Architectures
III-D1 Latency
LEO satellites experience non-negligible propagation and routing delays, on average around 25 ms [155, 71] of a round-trip time (RTT). For instance, a satellite operating at an altitude of approximately 550 km has a theoretical RTT propagation delay of about 3.6 ms; however, practical RTTs are significantly higher, ranging between 30 and 50 ms, due to routing, ground station access, and network management overheads. Notably, in our experimental evaluation [67], we revealed periodic high-magnitude RTT spikes of up to 140 ms, at regular intervals of approximately 13 s. This periodic behavior is indicative of constellation-specific control and resource management mechanisms, such as inter-beam or inter-satellite handovers, which require additional signaling for path reconfiguration and session re-initialization. By contrast, HAPS platforms operating in the stratosphere at altitudes of 18–25 km have nearly negligible propagation delays, and the resulting RTTs is typically comparable to that of terrestrial 5G systems, consistently below 10 ms. This makes HAPS particularly attractive for latency-sensitive services. Finally, UAV swarms operate at very low altitudes (e.g., a few hundred meters), and achieve sub-millisecond propagation delays. However, in practice, the end-to-end latency is often dominated by multi-hop processing and coordination overhead within the swarm’s mesh network [152, 143].
III-D2 CAPEX/OPEX
Global LEO constellations require extreme upfront capital investment. Public estimates for Amazon’s Project Kuiper and SpaceX’s Starlink indicate a total CAPEX of approximately 10 billion USD to reach full operational capability [168, 132]. According to [236], the average CAPEX per Starlink satellite is approximately 548,000 USD, for both satellite manufacturing and launch costs. Inter-satellite links (ISLs) in second-generation satellites introduce an additional cost of several hundred thousand USD per link. Regarding OPEX, costs are primarily dominated by ground stations, which require an average annual cost of approximately 215,000 USD on average, vs. around 20,000 USD for each satellite. Furthermore, the relatively short operational lifetime of LEO satellites (approximately five years) necessitates frequent satellite replacement, thereby increasing long-term costs. In comparison, the CAPEX of HAPS is significantly lower. For example, the manufacturing cost of a Airbus Zephyr platform is estimated at approximately 4 million USD [212], a number that has decreased significantly as stratospheric technology have matured. Unlike LEO satellites, HAPS platforms are typically designed to operate for around of 20 years, so that initial investments can be amortized over a much longer period. In terms of OPEX, costs are mainly due to maintenance and logistics, reaching up to 15% of the manufacturing cost annually. However, energy costs remain minimal as HAPS platforms utilize high-efficiency solar panels for daytime flight and onboard batteries for nighttime operations, resulting in an annual energy cost of only 30,000 USD per platform. Finally, the individual CAPEX of large civilian heavy-lift UAVs is relatively low, typically on the order of a few thousand USD, while OPEX is dominated by energy consumption and maintenance, in particular battery replacement and charging logistics.
III-D3 Scalability
While large-scale LEO satellite constellations are designed to provide ubiquitous coverage on Earth, the network capacity must be shared among a large number of UEs, which can saturate the available resources and complicate channel access [181]. In contrast, HAPS platforms only operate at a regional level, but provide high-capacity quasi-stationary connectivity via aggressive frequency reuse. Finally, UAV swarms provide extreme localized (though high-speed) coverage in hot-spot areas (e.g., for temporary events or emergency scenarios), even though they do not scale efficiently over large geographic regions.
III-D4 Operational risk
LEO constellations face both short- and long-term operational risks, including launch failures, orbital congestion, and the increasing threat of space debris and Kessler Syndrome effects. HAPS operations are primarily challenged by stratospheric wind dynamics and seasonal power constraints, particularly during the winter solstice at latitudes above approximately 50, where the shorter daylight duration may limit battery recharging capability for overnight flight. UAV swarms face the most stringent regulatory and legal constraints on the maximum flight altitude, airspace access, and spectrum usage, especially for Beyond Visual Line-of-Sight (BVLOS) flight operations over populated areas, which typically require case-by-case authorization and continuous human supervision. Moreover, UAVs are vulnerable to adverse weather conditions, such as strong winds and precipitation.
III-E Cross-Platform Energy Modeling for HAPS
A single energy-model framework that separates (i) solar energy harvesting, (ii) storage restrictions, and (iii) platform/payload power budgets is necessary for a consistent comparison of HAPS types (such as fixed-wing solar aircraft and aerostatic/airship systems). The reported energy efficiency and endurance findings are not directly comparable across research due to the lack of established measures and assumptions.
Solar harvesting is commonly represented as a function of solar irradiance $G(t)$ (W/m2), conversion efficiency $\eta_{\mathrm{PV}}$ (%), and effective photovoltaic area $A_{\mathrm{PV}}$ (m2). The harvested power can be expressed as
$P_{\mathrm{harv}}(t)=\eta_{\mathrm{PV}}\,A_{\mathrm{PV}}\,G(t)$ (1)
using temporally integrated daily gathered energy. To present harvested energy in kWh/day and facilitate cross-platform normalization, fixed-wing solar HAPS experiments explicitly parameterize these values under latitude- and season-dependent irradiance circumstances [29]. Similar irradiance-driven formulations are used in system-level HAPS energy studies, confirming their applicability as a platform-independent reference model [15].
Storage limits are characterized using the storage capacity $E_{\mathrm{sto}}$ (kWh), the usable fraction (depth-of-discharge and round-trip constraints), and the equivalent sustainment time under a defined load. The storage sustainment time is given by
$T_{\mathrm{sto}}=\frac{E_{\mathrm{use}}}{P_{\mathrm{tot}}}$ (2)
where $P_{\mathrm{tot}}$ denotes the total platform power demand. Studies on the viability and power budget of high-altitude airships show how storage sizing relates to propulsion or station-keeping requirements and payload loads. These works also provide guidance for consistent reporting of storage capacity and endurance across platform types [50, 46].
Payload power budgets are most comparable when decomposed into communications payload power $P_{\mathrm{pl}}$, propulsion or station-keeping power $P_{\mathrm{prop}}$, and auxiliary/avionics power $P_{\mathrm{aux}}$. The total platform power consumption can therefore be written as
$P_{\mathrm{tot}}=P_{\mathrm{pl}}+P_{\mathrm{prop}}+P_{\mathrm{aux}}$ (3)
This decomposition enables comparison between propulsion-dominant fixed-wing platforms and station-keeping-dominant aerostatic systems. In particular, the payload power share $\left(P_{\mathrm{pl}}/P_{\mathrm{tot}}\right)$ can be used as a normalized efficiency indicator [29, 46].
To ensure fair interpretation for TN–NTN integration research, these platform-level metrics should be reported alongside altitude, endurance targets, and irradiance assumptions (latitude and season). Furthermore, linking platform power budgets to network-level effects—such as reducing terrestrial densification requirements or supporting traffic spikes—strengthens the relevance of energy comparisons in integrated architectures [121]. Consequently, a transparent and repeatable basis for cross-platform evaluation can be achieved by presenting solar harvesting characteristics, storage capacity and sustainment, and complete power budget breakdowns within a unified comparison table.
III-F HAPS Market Overview
HAPS offer a cost-effective complement to terrestrial and satellite networks in NTN architectures, particularly for rural connectivity and emergency services [182]. This subsection addresses economic viability through cost breakdowns, sustainable business models, and market analysis of key providers. As mentioned above, the CAPEX/OPEX comparison between LEO satellites and HAPS is generally more favorable for HAPS, and the overall environmental burden is typically lower. Meanwhile, sustainable models for HAPS operators emphasize multi-layer integration with satellites and TN infrastructure, industry partnership, and green energy to achieve long-term viability [125]. Reliance on solar power supports UN Sustainable Development Goals 7 and 13, reducing carbon emissions and enabling eco-efficient operations in remote, disaster, and maritime scenarios. Collaborative models, such as those between Airbus, NTT DOCOMO, and SKY Perfect JSAT [166] or SoftBank and Sceye [199], focus on direct-to-device services via HAPS-NTN hybrids, targeting air/sea connectivity and disaster response. These ensure sustainability through spectrum sharing, regulatory alignment, and revenue from government/defense contracts (e.g., surveillance, environmental monitoring).
While the global HAPS market is clearly evolving and expanding across a broad range of applications, it remains nascent and fragmented. The limited availability of long-term historical data makes market projections highly sensitive to assumptions regarding adoption rates and regulatory developments. According to MarkNtel Advisors, the HAPS market is projected to grow from \$99 million in 2024 to \$240 million by 2030, driven by a diverse set of industrial applications [151]. Published in March 2025, this report focuses primarily on core HAPS platforms, excluding ancillary services, ground infrastructure, and broader ecosystem components from its scope. In contrast, Global Insight Services estimates a significantly larger market expansion—from \$1.7 billion in 2025 to \$3.3 billion by 2035—anticipating growth supported by advances in materials science, solar power technologies, and regulatory facilitation [79]. This assessment adopts a broader ecosystem perspective, encompassing not only the platforms themselves but also associated products, subsystems, and supporting equipment. Despite differences in scope and valuation methodology, both reports converge on the conclusion that the communications segment constitutes the dominant share of the HAPS market (approximately 35–45%), followed by surveillance and navigation applications, with the Asia-Pacific region emerging as a leading growth area.
IV HAPS Use Cases